a mother is 46 years old and her son 21 years old in the year 1997.In what year was the mother six times as old as son?
Answer by ankor@dixie-net.com(22740)

a mother is 46 years old and her son 21 years old in the year 1997.
In what year was the mother six times as old as son?
:
Let y = no. yrs ago Mom was 6 times as old as the son
:
46 - y = 6(21 - y)
46 - y = 126 - 6y
-y + 6y = 126 - 46
5y = 80
y = 80/5
y = 16 yrs ago
:
1997 - 16 = 1981 Mom was 6 times son
:
:
Confirm this by finding how old each was 16 yr ago
46 - 16 = 30
21 - 16 = 5
6 * 5 = 30